# Sound Clouds: Exploring ambient intelligence in public spaces to elicit deep human experience of awe, wonder, and beauty

### v1

@@ 1 Introduction
Ambient intelligence (AmI), as a subset of artificial intelligence, has started to play an increasingly inseparable role in our daily technological encounters. While there are varied definitions of AmI, one widely adopted definition from Aarts for ambient intelligence is “developing technology that will increasingly make our everyday environment sensitive and responsive to our presence” Aarts and Encarnação (2006). AmI generally possesses the characteristics of being Sensitive (S), Responsive (R), Adaptive (A), Transparent (T), Ubiquitous (U), and Intelligent (I) Cook et al. (2009). While most of the ambient intelligence serves a pragmatic purpose, such as a) the Nest thermostat Hernandez et al. (2014) for achieving domestic energy-saving, b) ambient influence for promoting positive behavior change like taking the stairs Rogers et al. (2010), and c) LAICA for enhancing public security Cucchiara et al. (2005), etc., there has been little exploration in other kinds of human experience with ambient intelligence. Therefore, we ask the following overarching research question: How can we design interactions with ambient artificial intelligence in public spaces to elicit deep human feelings of awe, wonder, and beauty?
@@ 2.1 The roles of AI and ML
We used computer vision (CV) for tracking spheres using a fine-tuned YOLO12n model  [^3] YOLO12n https://docs.ultralytics.com/models/yolo12/. The music generation is rule-based through Max/MSP  [^4] Max/MSP Wikipedia https://en.wikipedia.org/wiki/Max_(software) software. Although the CV model tracked all the spheres, we used the parameters of two spheres for music generation. We looped an ocean-like sound as the background, depicting the ethereal, mysterious atmosphere. We divided the video feed space into nine zones and used three key scripts for sound generation (see the description below and Figure 3, Rigt). Due to the fish’s-eye view of the GoPro wide-angle lens, the vision is lightly distorted (see Figure 3, Left). The key scripts for interaction are:
@@ 3 Sound Clouds and Humanity
While the ambient intelligence usually serves a pragmatic purpose, Sound Clouds elicits deep human experience of awe, wonder, and beauty. Awe, as a specific self-transcendent emotion, is characterized by “the perception of being in the presence of something vast that the individual does not immediately understand” Stellar et al. (2017). As a rarely-experienced emotion, experiencing awe has been found to prompt feelings of interconnectedness Shiota et al. (2007), motivate commitment to social collectives Keltner and Haidt (2003), and improve perception of well-being Rudd et al. (2012). Thus, experiencing awe can contribute to the exploration of humanity aspects of AI. Participants shared their feelings of awe during the post-interaction interview, including:
“…the ambience and the whole vibe of the place was very peaceful and calming
@@
Following the research thread of The Aware Home Abowd et al. (2002), an AmI research space supporting the elderly’s dignified aging, and Smart Kitchen Bell and Kaye (2002)–which prioritizes “experience, affect, and desire” but not just efficiency, our exhibit serves as a provocation for future AmI systems that enrich–not replace–humanity.
@@ 4 Author Bibliography
  • Chengzhi Zhang is a PhD student at Georgia Tech with a research focus on AI literacy Zhang and Magerko (2025) and the intersection of AI and UX Zhang et al. (2023).
  • Dashiel Carrera is a Visiting PhD Researcher at Columbia University, a novelist, and a sound media artist whose research and art practice explore AI’s impact on the Arts Carrera et al. (2025).
@@
  • Jasmine Kaur is a graduate student at Georgia Tech with a research focus on human-computer interaction and speculative design with a special emphasis on creative AI technologies  Kaur et al. (2025).
  • Jisu Kim is an M.S. student and Graduate Research Assistant at Georgia Tech with a research focus on responsible AI Lee et al. (2024) and human-AI collaboration, specifically how AI can augment–rather than replace–human values.

### v2

@@ 1 Introduction
Ambient intelligence (AmI), as a subset of artificial intelligence, has started to play an increasingly inseparable role in our daily technological encounters. While there are varied definitions of AmI, one widely adopted definition from  Aarts and Encarnação (2006) for ambient intelligence is “developing technology that will increasingly make our everyday environment sensitive and responsive to our presence”. AmI generally possesses the characteristics of being Sensitive (S), Responsive (R), Adaptive (A), Transparent (T), Ubiquitous (U), and Intelligent (I) (Cook et al., 2009). While most of the ambient intelligence serves a pragmatic purpose, such as a) the Nest thermostat (Hernandez et al., 2014) for achieving domestic energy-saving, b) ambient influence for promoting positive behavior change like taking the stairs (Rogers et al., 2010), and c) LAICA for enhancing public security (Cucchiara et al., 2005), etc., there has been little exploration in other kinds of human experience with ambient intelligence. Therefore, we ask the following overarching research question: How can we design interactions with ambient artificial intelligence in public spaces to elicit deep human feelings of awe, wonder, and beauty?
@@ 2.1 The roles of AI and ML
We used computer vision (CV) for tracking spheres using a fine-tuned YOLO12n model  [^3] YOLO12n https://docs.ultralytics.com/models/yolo12/. The music generation is rule-based through Max/MSP  [^4] Max/MSP Wikipedia https://en.wikipedia.org/wiki/Max_(software) software. Although the CV model tracked all the spheres, we used the parameters of two spheres for music generation. We looped an ocean-like sound as the background, depicting the ethereal, mysterious atmosphere. We divided the video feed space into nine zones and used three key scripts for sound generation (see the description below and Figure 3, Rigt). Due to the fish-eye’s view of the GoPro wide-angle lens, the vision is lightly distorted (see Figure 3, Left). The key scripts for interaction are:
@@ 3 Sound Clouds and Humanity
While the ambient intelligence usually serves a pragmatic purpose, Sound Clouds elicits deep human experience of awe, wonder, and beauty. Awe, as a specific self-transcendent emotion, is characterized by “the perception of being in the presence of something vast that the individual does not immediately understand” (Stellar et al., 2017). As a rarely-experienced emotion, experiencing awe has been found to prompt feelings of interconnectedness (Shiota et al., 2007), motivate commitment to social collectives (Keltner and Haidt, 2003), and improve perception of well-being (Rudd et al., 2012). Thus, experiencing awe can contribute to the exploration of humanity aspects of AI. Participants shared their feelings of awe during the post-interaction interview, including:
“…the ambiance and the whole vibe of the place was very peaceful and calming
@@
Following the research thread of The Aware Home (Abowd et al., 2002), an AmI research space supporting the elderly’s dignified aging, and Smart Kitchen (Bell and Kaye, 2002)–which prioritizes “experience, affect, and desire” but not just efficiency, our exhibit serves as a provocation for future AmI systems that enrich–not replace–humanity.
@@ 4 Author Bibliography
  • Chengzhi Zhang is a PhD student at Georgia Tech with a research focus on AI literacy (Zhang and Magerko, 2025) and the intersection of AI and UX (Zhang et al., 2023).
  • Dashiel Carrera is a Visiting PhD Researcher at Columbia University, a novelist, and a sound media artist whose research and art practice explore AI’s impact on the Arts (Carrera et al., 2025).
@@
  • Jasmine Kaur is a graduate student at Georgia Tech with a research focus on human-computer interaction and speculative design with a special emphasis on creative AI technologies (Kaur et al., 2025).
  • Jisu Kim is an M.S. student and Graduate Research Assistant at Georgia Tech with a research focus on responsible AI (Lee et al., 2024) and human-AI collaboration, specifically how AI can augment–rather than replace–human values.